# Characterization of fractional Chern insulator quasiparticles in moiré transition metal dichalcogenides

### v1

@@ III.1 Model of tMoTe2
[…]
@@
in the layer basis, with $t$ and $b$ denoting the top and bottom layer, respectively. The Hamiltonian $H_{0,K}({\bm{r}})$ differs from that in Ref. [87] by a particle-hole transformation. Here $h_{l}({\bm{r}})=-\hbar(+i\nabla-{\bm{K}}_{l})^{2}/(2m^{*})$ describes the quadratic valence band edge of monolayer MoTe2, where $m^{*}=0.6m_{e}$ is the effective mass and ${\bm{K}}_{l}$ is the $K$ point of layer $l$. Each layer feels a moiré potential $\Delta_{l}({\bm{r}})=2v\sum_{j=1,3,5}\cos({\bm{G}}_{j}\cdot{\bm{r}}+l\phi)$, where ${\bm{G}}_{j}=\frac{4\pi}{\sqrt{3}a_{M}}\left(\cos\left(\frac{\pi}{3}(j-1)\right),\sin\left(\frac{\pi}{3}(j-1)\right)\right)$ are the shortest moiré reciprocal lattice vectors, and $l$ is $+1$ ($-1$) for the bottom (top) layer. The interlayer tunneling $\Delta_{T}({\bm{r}})=w(1+e^{-i{\bm{G}}_{2}\cdot{\bm{r}}}+e^{-i{\bm{G}}_{3}\cdot{\bm{r}}})$. Both the intralayer moiré potential and the interlayer tunneling are approximated by first harmonic expansion. Throughout our work, we adopt $(v,\phi,w)=(20.8\ {\rm meV},+107.7^{\circ},-23.8\ {\rm meV})$ [37], which were obtained by recent first-principles calculations for $\theta$ near $4^{\circ}$. The single-particle Hamiltonian in the $K^{\prime}$ valley can be determined by time reversal conjugate of $H_{0,K}({\bm{r}})$.
@@
$H_{0,K}$ $=$ $-\sum_{\bm{k}}\Bigg{\{}\sum_{l=t,b}h_{l}(-{\bm{k}}-{\bm{K}}_{l})c^{\dagger}_{K,l,{\bm{k}}}c_{K,l,{\bm{k}}}+v\sum_{l=t,b}\sum_{j=1,3,5}\left(e^{il\phi}c^{\dagger}_{K,l,{\bm{k}}+{\bm{G}}_{j}}c_{K,l,{\bm{k}}}+{\rm h.c.}\right)$ (8)
$+$ $w\left(c^{\dagger}_{K,b,{\bm{k}}}c_{K,t,{\bm{k}}}+c^{\dagger}_{K,b,{\bm{k}}+{\bm{G}}_{2}}c_{K,t,{\bm{k}}}+c^{\dagger}_{K,b,{\bm{k}}+{\bm{G}}_{3}}c_{K,t,{\bm{k}}}+{\rm h.c.}\right)\Bigg{\}},$
where $c^{\dagger}_{K,l,{\bm{k}}}$ creates a hole with wave vector ${\bm{k}}$ in valley $K$ and layer $l$. The monolayer Hamiltonian is $h_{l}(-{\bm{k}}-{\bm{K}}_{l})=-\frac{\hbar^{2}|-{\bm{k}}-{\bm{K}}_{l}|^{2}}{2m^{*}}$. For each point ${\bm{k}}_{0}$ in the MBZ of the $K$ valley, we let ${\bm{k}}={\bm{k}}_{0}+m_{1}{\bm{G}}_{1}+m_{2}{\bm{G}}_{2}$ in Eq. (8). With truncations $m_{1},m_{2}=-M,\cdots M$ (we choose $M=7$ in our calculations), $H_{0,K}$ can be constructed as a matrix of dimension $2(2M+1)^{2}$ under the plane-wave basis $|K,l,{\bm{k}}\rangle\equiv c^{\dagger}_{K,l,{\bm{k}}}|{\rm vac}\rangle$. The eigenvalues $E_{K,n,{\bm{k}}_{0}}$ of this matrix give the hole band energies at ${\bm{k}}_{0}$, with $n$ the band index. The band eigenvectors can be expressed as $|K,n,{\bm{k}}_{0}\rangle=\sum_{l,m_{1},m_{2}}u_{l,m_{1},m_{2}}^{K,n,{\bm{k}}_{0}}|K,l,{\bm{k}}_{0}+m_{1}{\bm{G}}_{1}+m_{2}{\bm{G}}_{2}\rangle$. In real space, the corresponding Bloch state is a layer spinor
@@
where all wave vectors are restricted in the MBZ, and $\gamma_{\eta,n,{\bm{k}}}^{\dagger}$ creates a hole with wave vector ${\bm{k}}$ in band $n$ of valley $\eta$, i.e., $|\eta,n,{\bm{k}}\rangle=\gamma_{\eta,n,{\bm{k}}}^{\dagger}|{\rm vac}\rangle$. We have assumed that the impurity potential does not flip spins. As a result of spin-valley locking in tMoTe2, the impurity potential does not couple the two valleys and the number of holes in each valley is conserved. The interaction matrix element $V^{(\eta\eta^{\prime})}_{\{{\bm{k}}_{i}\}\{n_{i}\}}=\frac{1}{2}\delta^{\prime}_{{\bm{k}}_{1}+{\bm{k}}_{2},{\bm{k}}_{3}+{\bm{k}}_{4}}\sum_{\bm{G}}V({\bm{k}}_{1}-{\bm{k}}_{4}+{\bm{G}})M_{\eta,n_{1},n_{4}}({\bm{k}}_{1},{\bm{k}}_{4}-{\bm{G}})M_{\eta^{\prime},n_{2},n_{3}}({\bm{k}}_{2},{\bm{k}}_{3}+{\bm{G}}+\delta{\bm{G}})$, where $\delta^{\prime}$ is the 2D periodic Kronecker delta function with period of primitive moiré reciprocal lattice vectors, $V({\bm{q}})$ is the Fourier transform of the interaction potential of holes, ${\bm{G}}$ is a reciprocal lattice vector, $M_{\eta,n,n^{\prime}}({\bm{k}},{\bm{k}}^{\prime})\equiv\sum_{l,m_{1},m_{2}}[u_{l,m_{1},m_{2}}^{\eta,n,{\bm{k}}}]^{*}u_{l,m_{1},m_{2}}^{\eta,n^{\prime},{\bm{k}}^{\prime}}$, and $\delta{\bm{G}}={\bm{k}}_{1}+{\bm{k}}_{2}-{\bm{k}}_{3}-{\bm{k}}_{4}$. Here we adopt the gauge that $\psi_{\eta,n,{\bm{k}}_{0}+\bm{G}}({\bm{r}})=\psi_{\eta,n,{\bm{k}}_{0}}({\bm{r}})$ for any reciprocal lattice vector $\bm{G}$. Therefore, $u_{l,m_{1},m_{2}}^{\eta,n,{\bm{k}}_{0}+\bm{G}^{\prime}}=u_{l,m_{1}+m_{1}^{\prime},m_{2}+m_{2}^{\prime}}^{\eta,n,{\bm{k}}_{0}}$, where $\bm{G}^{\prime}=m_{1}^{\prime}\bm{G}_{1}+m_{2}^{\prime}\bm{G}_{2}$. In the following, we still choose $V({\bm{q}})=\frac{e^{2}}{4\pi\epsilon\epsilon_{0}}\frac{1}{\mathcal{A}}\frac{2\pi\tanh(|{\bm{q}}|d)}{|{\bm{q}}|}$, as in the FQH case. The matrix element of the impurity potential is $U^{\eta}_{\{{\bm{k}}_{i}\}\{n_{i}\}}=\sum_{\bm{G}}U({\bm{k}}_{1}-{\bm{k}}_{2}+{\bm{G}})M_{\eta,n_{1},n_{2}}({\bm{k}}_{1},{\bm{k}}_{2}-{\bm{G}})$, where $U({\bm{q}})$ is the Fourier transform of the impurity potential. For numerical efficiency, one has to project Eq. (10) to a finite set of bands. The experimental observations of valley polarized FCIs at hole filling $\nu_{h}=2/3$ have been supported by single-band ED that only keeps the top valence band in a single valley [37, 38]. ED calculations that keep both valleys and two bands per valley also indicate that the $\nu_{h}=2/3$ FCIs still survive in a small region of the $(\theta,\epsilon)$ parameter space [40].
@@ III.2 Density of holes in FCI ground states
[…]
Figure 3: (a) The real-space density of holes for the $\nu_{h}=2/3$ FCI ground state in tMoTe2. The system size is $N=24,N_{s}=36$. The white dots indicate the $\mathcal{R}_{M}^{M}$ positions. In (b), we show the momentum-space occupation for the system in (a). We set $\theta=3.7^{\circ}$, $d=10\ {\rm nm}$ and $\epsilon=10$.
Unlike the FQH case, the real-space density of moiré FCIs is not uniform, as shown in Fig. 3(a). $\rho({\bm{r}})$ is strongly peaked near the $\mathcal{R}_{M}^{X}$ and $\mathcal{R}_{X}^{M}$ positions, but almost vanishes near the $\mathcal{R}_{M}^{M}$ points. As the hole occupation in the momentum space is approximately uniform [Fig. 3(b)], this non-uniform $\rho({\bm{r}})$ is determined by the real-space distribution of single-particle band wave functions. As shown in Ref. [90], the holes in the top valence band indeed tend to be confined near the $\mathcal{R}_{M}^{X}$ and $\mathcal{R}_{X}^{M}$ positions when the first two valence bands carry Chern number $\mathcal{C}=\pm 1$, which holds for the parameters chosen by us. In this case, the first two valence bands in the $K$ valley can be approximated by the Haldane model on an effective honeycomb lattice formed by $\mathcal{R}_{M}^{X}$ and $\mathcal{R}_{X}^{M}$ sites. We emphasize that the density non-uniformity only occurs in the length scale of one moiré unit cell for the FCI ground states. After summing the density within a moiré unit cell, we get the same total particle number $n_{\rm uc}=|\nu|/S_{\rm uc}$ for all unit cells, where $S_{\rm uc}$ is the area of a moiré unit cell.
@@ IV Quasiparticles in FCIs
Now we study quasiparticles for the $\nu_{h}=2/3$ moiré FCIs in tMoTe2. Note that the particle in “quasiparticle” actually means hole here. To generate $N_{\rm qp}$ quasiparticles, we require the number of moiré unit cells $N_{s}$ and the number of holes $N$ to satisfy $N_{s}=\frac{1}{2}(3N-N_{\rm qp})$. It is natural to use a delta impurity potential $U_{\rm imp}({\bm{w}})$ of strength $U_{0}$ at position ${\bm{w}}$ that is attractive for holes to localize a quasiparticle at ${\bm{w}}$. The Fourier transform of the impurity potential is $U({\bm{q}})=\frac{U_{0}}{N_{s}}e^{-i{\bm{q}}\cdot{\bm{w}}}$. In this section, we choose $d=10\ {\rm nm}$ and $\epsilon=10$, and project the Hamiltonian to the top valence band of the $K$ valley. We will consider other parameter values and include more bands in Appendix C and D.
@@
Figure 5: The low-energy energy spectra in presence of (a) a single delocalized quasiparticle and (b) two delocalized quasiparticles. The system sizes are $N=23,N_{s}=34$ in (a) and $N=24,N_{s}=35$ in (b). $Q$ is a proper folding of the two-dimensional momentum ${\bm{Q}}$. The definitions of $E_{\rm 1qp}$ and $V_{\rm 2qp}$ are given in the text. The subspace of delocalized quasiparticles are highlighted by red. This subspace contains one level per momentum sector in (a), and six levels per momentum sector in (b). The gray levels are out of this subspace. We set $\theta=3.7^{\circ}$, $d=10\ {\rm nm}$ and $\epsilon=10$.

### v3

@@ III.1 Model of tMoTe2
The moiré superlattice consists of two MoTe2 monolayers of lattice constant $a=3.52\ \mathring$$\mathrm${A}}, with a twist angle $\theta$ in between. At small twist angles, the lattice constant of the moiré pattern is $a_{M}\approx a/\theta\gg a$. Accordingly, the monolayer Brillouin zone is folded into the much smaller moiré Brillouin zone (MBZ). There are three high-symmetry positions in each moiré unit cell: $\mathcal{R}_{M}^{M}$, $\mathcal{R}_{M}^{X}$ and $\mathcal{R}_{X}^{M}$, where $M$ and $X$ represent metal and chalcogen atoms, respectively. $\mathcal{R}_{\alpha}^{\beta}$ stands for the local configuration in which $\alpha$ atom in the bottom layer is vertically aligned with the $\beta$ atom in the top layer. We choose the primitive moiré lattice vectors as ${\bm{a}}_{1}=(\frac{\sqrt{3}}{2},\frac{1}{2})a_{M}$ and ${\bm{a}}_{2}=(-\frac{\sqrt{3}}{2},\frac{1}{2})a_{M}$. In the reciprocal space, the $K$ and $K^{\prime}$ valleys are separated by a large momentum difference at small twist angles $\theta$, such that they can be considered independently. Moreover, spin is effectively locked to valley due to the strong spin-orbit coupling near the two valleys.
@@
in the layer basis, with $t$ and $b$ denoting the top and bottom layer, respectively. The Hamiltonian $H_{0,K}({\bm{r}})$ differs from that in Ref. [92] by a particle-hole transformation. Here
$h_{l}({\bm{r}})=-\frac{\hbar}{2m^{*}}(+i\nabla-{\bm{K}}_{l})^{2}$ (12)
describes the quadratic valence band edge of monolayer MoTe2, where $m^{*}=0.6m_{e}$ is the effective mass and ${\bm{K}}_{l}$ is the $K$ point of layer $l$. Each layer feels a moiré potential
$\Delta_{l}({\bm{r}})=2v\sum_{j=1,3,5}\cos({\bm{G}}_{j}\cdot{\bm{r}}+l\phi),$ (13)
where ${\bm{G}}_{j}=\frac{4\pi}{\sqrt{3}a_{M}}\left(\cos\left(\frac{\pi}{3}(j-1)\right),\sin\left(\frac{\pi}{3}(j-1)\right)\right)$ are the shortest moiré reciprocal lattice vectors, and $l$ is +1+1 (1-1) for the bottom (top) layer. The interlayer tunneling is
$\Delta_{T}({\bm{r}})=w(1+e^{-i{\bm{G}}_{2}\cdot{\bm{r}}}+e^{-i{\bm{G}}_{3}\cdot{\bm{r}}}).$ (14)
Both the intralayer moiré potential and the interlayer tunneling are approximated by first harmonic expansion. Throughout our work, we adopt $(v,\phi,w)=(20.8\ {\rm meV},+107.7^{\circ},-23.8\ {\rm meV})$ [37], which were obtained by recent first-principles calculations for $\theta$ near $4^{\circ}$. The single-particle Hamiltonian in the $K^{\prime}$ valley can be determined by time reversal conjugate of $H_{0,K}({\bm{r}})$.
@@
$H_{0,K}$ $=$ $-\sum_{\bm{k}}\Bigg\{\sum_{l=t,b}h_{l}(-{\bm{k}}-{\bm{K}}_{l})c^{\dagger}_{K,l,{\bm{k}}}c_{K,l,{\bm{k}}}$ (15)
$+$ $v\sum_{l=t,b}\sum_{j=1,3,5}\left(e^{il\phi}c^{\dagger}_{K,l,{\bm{k}}+{\bm{G}}_{j}}c_{K,l,{\bm{k}}}+{\rm h.c.}\right)$
$+$ $w\Big(c^{\dagger}_{K,b,{\bm{k}}}c_{K,t,{\bm{k}}}+c^{\dagger}_{K,b,{\bm{k}}+{\bm{G}}_{2}}c_{K,t,{\bm{k}}}$
$+$ $c^{\dagger}_{K,b,{\bm{k}}+{\bm{G}}_{3}}c_{K,t,{\bm{k}}}+{\rm h.c.}\Big)\Bigg\},$
where $c^{\dagger}_{K,l,{\bm{k}}}$ creates a hole with wave vector $\bm{k}$} in valley $K$ and layer $l$. The monolayer Hamiltonian is $h_{l}(-{\bm{k}}-{\bm{K}}_{l})=-\frac{\hbar^{2}|-{\bm{k}}-{\bm{K}}_{l}|^{2}}{2m^{*}}$. For each point ${\bm{k}}_{0}$ in the MBZ of the $K$ valley, we let ${\bm{k}}={\bm{k}}_{0}+m_{1}{\bm{G}}_{1}+m_{2}{\bm{G}}_{2}$ in Eq. (15). With truncations $m_{1},m_{2}=-M,\cdots M$ (we choose $M=7$ in our calculations), $H_{0,K}$ can be constructed as a matrix of dimension $2(2M+1)^{2}$ under the plane-wave basis $|K,l,{\bm{k}}\rangle\equiv c^{\dagger}_{K,l,{\bm{k}}}|{\rm vac}\rangle$. The eigenvalues $E_{K,n,{\bm{k}}_{0}}$ of this matrix give the hole band energies at ${\bm{k}}_{0}$, with $n$ the band index. The band eigenvectors can be expressed as
$|K,n,{\bm{k}}_{0}\rangle=\sum_{l,m_{1},m_{2}}u_{l,m_{1},m_{2}}^{K,n,{\bm{k}}_{0}}|K,l,{\bm{k}}_{0}+m_{1}{\bm{G}}_{1}+m_{2}{\bm{G}}_{2}\rangle.$
In real space, the corresponding Bloch state is a layer spinor
@@
where all wave vectors are restricted in the MBZ, and $\gamma_{\eta,n,{\bm{k}}}^{\dagger}$ creates a hole with wave vector $\bm{k}$} in band $n$ of valley $\eta$, i.e., $|\eta,n,{\bm{k}}\rangle=\gamma_{\eta,n,{\bm{k}}}^{\dagger}|{\rm vac}\rangle$. We have assumed that the impurity potential does not flip spins. As a result of spin-valley locking in tMoTe2, the impurity potential does not couple the two valleys and the number of holes in each valley is conserved. Defining
$M_{\eta,n,n^{\prime}}({\bm{k}},{\bm{k}}^{\prime})\equiv\sum_{l,m_{1},m_{2}}[u_{l,m_{1},m_{2}}^{\eta,n,{\bm{k}}}]^{*}u_{l,m_{1},m_{2}}^{\eta,n^{\prime},{\bm{k}}^{\prime}},$ (19)
we can express the interaction matrix element as
$V^{(\eta\eta^{\prime})}_{\{{\bm{k}}_{i}\}\{n_{i}\}}=\frac{1}{2}\delta^{\prime}_{{\bm{k}}_{1}+{\bm{k}}_{2},{\bm{k}}_{3}+{\bm{k}}_{4}}\sum_{\bm{G}}V({\bm{k}}_{1}-{\bm{k}}_{4}+{\bm{G}})$
$\times M_{\eta,n_{1},n_{4}}({\bm{k}}_{1},{\bm{k}}_{4}-{\bm{G}})M_{\eta^{\prime},n_{2},n_{3}}({\bm{k}}_{2},{\bm{k}}_{3}+{\bm{G}}+\delta{\bm{G}}),$
where $\delta^{\prime}$ is the 2D periodic Kronecker delta function with period of primitive moiré reciprocal lattice vectors, $V({\bm{q}})$ is the Fourier transform of the interaction potential of holes, $\bm{G}$} is a reciprocal lattice vector, and $\delta{\bm{G}}={\bm{k}}_{1}+{\bm{k}}_{2}-{\bm{k}}_{3}-{\bm{k}}_{4}$. Here we adopt the gauge that $\psi_{\eta,n,{\bm{k}}_{0}+\bm{G}}({\bm{r}})=\psi_{\eta,n,{\bm{k}}_{0}}({\bm{r}})$ for any reciprocal lattice vector $\bm{G}$. Therefore, $u_{l,m_{1},m_{2}}^{\eta,n,{\bm{k}}_{0}+\bm{G}^{\prime}}=u_{l,m_{1}+m_{1}^{\prime},m_{2}+m_{2}^{\prime}}^{\eta,n,{\bm{k}}_{0}}$, where $\bm{G}^{\prime}=m_{1}^{\prime}\bm{G}_{1}+m_{2}^{\prime}\bm{G}_{2}$. In the following, we still choose $V({\bm{q}})=\frac{e^{2}}{4\pi\epsilon\epsilon_{0}}\frac{1}{\mathcal{A}}\frac{2\pi\tanh(|{\bm{q}}|d)}{|{\bm{q}}|}$, as in the FQH case. The matrix element of the impurity potential is
$U^{\eta}_{\{{\bm{k}}_{i}\}\{n_{i}\}}=\sum_{\bm{G}}U({\bm{k}}_{1}-{\bm{k}}_{2}+{\bm{G}})M_{\eta,n_{1},n_{2}}({\bm{k}}_{1},{\bm{k}}_{2}-{\bm{G}}),$
where $U({\bm{q}})$ is the Fourier transform of the impurity potential. For numerical efficiency, one has to project Eq. (18) to a finite set of bands. The experimental observations of valley polarized FCIs at hole filling $\nu_{h}=2/3$ have been supported by single-band ED that only keeps the top valence band in a single valley [37, 38]. ED calculations that keep both valleys and two bands per valley also indicate that the $\nu_{h}=2/3$ FCIs still survive in a small region of the $(\theta,\epsilon)$ parameter space [40].
@@ III.2 Density of holes in FCI ground states
where the expectation $\langle\cdot\rangle$ is over the ground state. The restriction ${\bm{k}}={\bm{k}}^{\prime}$ is imposed if the ground state has translation symmetry. We project Eq. (18) to the top valence band of valley $K$, and calculate the density averaged over the three $\nu_{h}=2/3$ ground states at twist angle $\theta=3.7^{\circ}$. In this case, the real-space density can be simplified to $\rho({\bm{r}})=\sum_{\bm{k}}n_{\bm{k}}|\psi_{{\bm{k}}}({\bm{r}})|^{2}$, where $n_{\bm{k}}$ is the momentum-space occupation of holes and $\psi_{{\bm{k}}}({\bm{r}})$ is the wavefunction of the single band. To make finite-size systems as isotropic as possible, we apply ED to samples satisfying tilted geometry [93, 94]. The sample details used in this work are given in Appendix A.
Figure 3: (a) The real-space density of holes for the $\nu_{h}=2/3$ FCI ground state in tMoTe2. The system size is $N=24,N_{s}=36$. The white dots indicate the $\mathcal{R}_{M}^{M}$ positions. In (b), we show the momentum-space occupation (averaged over the three FCI ground states) for the system in (a). We set $\theta=3.7^{\circ}$, $d=10\ {\rm nm}$ and $\epsilon=10$.
Unlike the FQH case, the real-space density of moiré FCIs is not uniform. In Fig. 3(a), we show the data averaged over three nearly degenerate FCI ground states. $\rho({\bm{r}})$ is strongly peaked near the $\mathcal{R}_{M}^{X}$ and $\mathcal{R}_{X}^{M}$ positions, but almost vanishes near the $\mathcal{R}_{M}^{M}$ points. As the hole occupation in the momentum space is approximately uniform [Fig. 3(b)], this non-uniform $\rho({\bm{r}})$ is determined by the real-space distribution of single-particle band wave functions. As shown in Ref. [95], the holes in the top valence band indeed tend to be confined near the $\mathcal{R}_{M}^{X}$ and $\mathcal{R}_{X}^{M}$ positions when the first two valence bands carry Chern number $\mathcal{C}=\pm 1$, which holds for the parameters chosen by us. In this case, the first two valence bands in the $K$ valley can be approximated by the Haldane model on an effective honeycomb lattice formed by $\mathcal{R}_{M}^{X}$ and $\mathcal{R}_{X}^{M}$ sites. We emphasize that the density non-uniformity only occurs in the length scale of one moiré unit cell for the FCI ground states. After summing the density within a moiré unit cell, we get the same total particle number $n_{\rm uc}=|\nu|/S_{\rm uc}$ for all unit cells, where $S_{\rm uc}$ is the area of a moiré unit cell.
@@ IV Quasiparticles in FCIs
Now we study quasiparticles for the $\nu_{h}=2/3$ moiré FCIs in tMoTe2. Note that the particle in “quasiparticle” actually means hole here. Like in the LLL, we require the number of moiré unit cells $N_{s}$ and the number of holes $N$ to satisfy $N_{s}=\frac{1}{2}(3N-N_{\rm qp})$ to generate $N_{\rm qp}$ quasiparticles. It is natural to use a delta impurity potential $U_{\rm imp}({\bm{w}})$ of strength $U_{0}$ at position $\bm{w}$} that is attractive for holes to localize a quasiparticle at $\bm{w}$}. The Fourier transform of the impurity potential is $U({\bm{q}})=\frac{U_{0}}{N_{s}}e^{-i{\bm{q}}\cdot{\bm{w}}}$. In this section, we choose $d=10\ {\rm nm}$ and $\epsilon=10$ unless otherwise stated, and project the Hamiltonian to the top valence band of the $K$ valley. We will consider other parameter values and include more bands in Appendix C and D.
@@
Figure 5: The low-energy energy spectra in presence of (a) a single delocalized quasiparticle and (b), (c) two delocalized quasiparticles. The system sizes are $N=23,N_{s}=34$ in (a) and $N=24,N_{s}=35$ in (b) and (c). $Q$ is a proper folding of the two-dimensional momentum $\bm{Q}$}. The definitions of $E_{\rm 1qp}$ and $V_{\rm 2qp}$ are given in the text. Unlike in Ref. [72], we do not shift the energies in (a) to make the dispersion near zero energy. The subspace of delocalized quasiparticles are highlighted by red. This subspace contains one level per momentum sector in (a), and six levels per momentum sector in (b) and (c). The gray levels are out of this subspace. We set $\theta=3.7^{\circ}$ and $\epsilon=10$. The sample-to-gate distance is chosen as $d=10\ {\rm nm}$ in (b) and $d=2\ {\rm nm}$ in (c).